OpenStudy (anonymous):

A triangle has side lengths of 18 cm, 80 cm, and 81 cm. Classify it as acute, obtuse, or right

OpenStudy (anonymous):

An acute triangle has all internal angles less than 90 degrees. An obtuse triangle has one internal angle greater than 90 degrees. A right triangle has one internal angle equal to exactly 90 degrees. Using Soh Cah Toa, can you work out what this triangle would be?

OpenStudy (anonymous):

we can assume that c = 81 in pyth theorem since it is the longest line a^2 + b^2 > c^2 obtuse angle a^2 + b^2 = c^2 right angle a^2 + b^2 < c^2 acute use a = 18 and b = 80